younggmoneyy:

Maria looks at the architectural plan of a four-walled room in which the walls meet each other at right angles. The length of one wall in the plan is 13 inches. The length of the diagonal of the floor of the room in the plan is approximately 15.26 inches. Is the room in the shape of a square? Explain how you determined your answer. Show all your work.

younggmoneyy:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @ZombieSlayer Here, Just apply the Pythagoras Theorem: Let, Unknown side = x Then, x² + 13² = 15.26² x² = 232.86 - 169 x = √63.87 x = 8 Here, Length of sides of the shape would be 8 & 13, so it is rectangle not a square as they are not equal Hope this helps! \(\color{#0cbb34}{\text{End of Quote}}\) thank you

ZombieSlayer:

No problemo!!

Jayden156:

yo mamaa

ZombieSlayer:

uhh